Problems 7–9 pertain to the information in the following two-way contingency table relating the opinion of a company’s employees on a proposed revision of the company pension plan and position in the company. Blue-collar White-collar Managerial For 0.335 0.160 0.055 Against 0.315 0.090 0.045

7. The probability that a randomly selected worker is in favor of the revision is about: (a) 0.65 (b) 0.45 (c) 0.10 (d) 0.38 (e) 0.55

8. The probability that a randomly selected worker is in favor of the revision, given that he is a white-collar worker, is about: (a) 0.16 (b) 0.29 (c) 0.64 (d) 0.47 (e) 0.72

9. The events F: the person selected is in favor of the revision and W: the person selected is a white-collar worker are: (a) independent because of P(F ∩ W) = P(F) · P(W). (b) independent because P(F ∪ W) = P(F) + P(W). (c) independent because P(F|W) = P(W). (d) dependent because P(F ∩ W) 6= P(F) · P(W). (e) dependent because P(F ∪ W) 6= P(F) + P(W).

Problems 10 and 11 pertain to the probability distribution of the number X of courses a randomly selected 2016 graduate of a certain university dropped in the course of his university studies. x 0 1 2 3 4 5 6 7 P(x) 0.06 0.09 0.18 0.26 0.23 0.13 0.04 0.01

10. The probability that a randomly selected 2016 graduate dropped at most two courses is about: (a) 0.15 (b) 0.33 (c) 0.18 (d) 0.67 (e) 0.85

11. The average number of courses dropped by 2016 graduates is about: (a) 3.62 (b) 2.13 (c) 3.50 (d) 3.11 (e) 4.00
Answer #1

Answer 10

P(At most 2 courses) = P(0 Courses) + P(1 Course) + P(2 Courses)

P(At most 2 courses) = 0.06 + 0.09 + 0.18

P(At most 2 courses) = 0.33 (Option B)

Answer 11

The average number of courses = Σx*P(x)

The average number of courses = 0*0.06 + 1*0.09 + 2*0.18 + 3*0.26 + 4*0.23 + 5*0.13 + 6*0.04 + 7*0.01

The average number of courses = 3.11 (Option D)

***Dear Student, We can answer sub-parts of one question per post. Please post sub-parts of other question separately***